A solution of Ca(OH)2 has a measured pH of 12.50.

What is the molar concentration of the Ca(OH)2 in the solution? What is the molar concentration of Ca(OH)2 if the solution is diluted so that the pH is 11.30?

Ca(OH)2 in the original solution = × 10 M

Ca(OH)2 in the diluted solution = × 10 M
Answer #1

1)

pH = -log [H3O+]

12.5 = -log [H3O+]

[H3O+] = 3.162*10^-13 M

use:

[OH-] = (1.0*10^-14)/[H3O+]

[OH-] = (1.0*10^-14)/(3.162*10^-13)

[OH-] = 3.162*10^-2 M

1 mol of Ca(OH)2 gives 2 mol of OH-

So,

[Ca(OH)2] = [OH-] / 2

= (3.162*10^-2)/2

= 0.01581 M

Answer: 0.01581 M

2)

pH = -log [H3O+]

11.30 = -log [H3O+]

[H3O+] = 5.012*10^-12 M

use:

[OH-] = (1.0*10^-14)/[H3O+]

[OH-] = (1.0*10^-14)/(5.012*10^-12)

[OH-] = 1.995*10^-3 M

1 mol of Ca(OH)2 gives 2 mol of OH-

So,

[Ca(OH)2] = [OH-] / 2

= (1.995*10^-3)/2

= 9.975*10^-4 M

Answer:9.975*10^-4 M